Jones dies and leaves £6874. He directs that the
money be divided among wife and 3 kids,so that
his wife receives 4 times as much as each child
Children are to get equal share.How much does
is left to the wife?

Answer by math_tutor2020(3816)


x = amount each child gets (in pounds)
4x = amount the wife gets

The 3 kids get x pounds each, so far that is x+x+x = 3x pounds.
Add on the wife's share to get 3x+4x = 7x
This total is set equal to £6874 and we solve for x.
Divide both sides by 7 to undo the "7 times" applied to x.

7x = 6874
7x/7 = 6874/7
x = 982

Each child gets £982
The wife gets 4x = 4*982 = 3928 pounds

Check:
child1+child2+child3+wife = 982+982+982+3928 = 6874
The answer is confirmed

Answer: £3928